When an aqueous solution containing gold (III) salt is electrolyzed, metallic gold is deposited at the cathode and oxygen gas is generated at the anode. If 9.26 g of Au is deposited at the cathode, calculate the volume (in liters) of O2 generated at 23°C and 747 mmHg.

1
Expert's answer
2021-12-24T04:05:33-0500

Molar mass of Au=196.97

9.26/196.97=0.047

= 0.047×295×747= 10357.16

= 10.3572L